OpenStudy (anonymous):

A rectangle has a length of 10 inches and a perimeter of 48 inches. Which equation could be used to find the width of the rectangle? A. 48 = 20 – 2w B. 48 = 10w C. 48 = 20 + 2w D. 48 = 40w

OpenStudy (anonymous):

@Skiller8860 plz help me

OpenStudy (anonymous):

The formula for the perimeter of your rectangle is 2L+2W=48 We know your length so let's plug that in.. 2(10)+2W=48

OpenStudy (anonymous):

What does 2 * 10 give you?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Correct, so we now simplify for W. 20+2W==48 ___________________ How do we get rid of the 20 on the side of 2W?

OpenStudy (anonymous):

Do you know how?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

We subtract 20 from both sides of the equation. Meaning we do this.. 20+2W=48 -20 -20 ------------ 2W= ... What does 48-20 give you?

OpenStudy (phi):

it would be good to be able to solve for w, but the question only asks for the equation

OpenStudy (anonymous):

28

OpenStudy (anonymous):

Correct so you have left over 2W=28 Now we get rid of the 2 beside the variable W. Opposite of multiplication is what?

OpenStudy (anonymous):

division

OpenStudy (anonymous):

Correct, so we will have to divide both sides by 2 to get W (width) by it's self. So 2 divided by 2 cancels that out. What we do to one side, we must to do the other, Now what is 28 divided by 2?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

So the final width would be 14. Now your answer to solve that problem would be C. 48 = 20 + 2w Thought I'd walk you through the steps so I could thoroughly explain it :).

OpenStudy (anonymous):

lol thx so much

OpenStudy (anonymous):

Anytime, glad I could help you.

OpenStudy (mrcoolguy):

–52 – (–27)

OpenStudy (mrcoolguy):

2 years ago